# Data-Driven Reachability with Scenario Optimization and the Holdout Method

### v1

@@ I Introduction
Reachability analysis plays an integral role in analyzing the safety of systems by determining the set of states a system could transition to in a finite time horizon [1]. Many set-based techniques [2, 3, 4, 5] have been developed to compute either under- or over-approximated reachable sets to capture all possible trajectories and present guarantees on state-space safety. However, when there is limited information on a system’s dynamics, through only simulation or experimentation, we must estimate reachable sets and derive probabilistic guarantees of correctness directly from data. To provide probability measures for learned reachable sets, many approaches require a minimum number of samples [6, 7, 8, 9] or assumptions on the form of system dynamics [10, 11, 12], and are often computationally intensive [13, 14]. We advocate for the use of a general purpose method, the holdout method, that circumvents these limitations and improves upon the sample complexity and sharpness of existing probabilistic reachability bounds.
The holdout method, or cross-validation, is one of the simplest and most widely used methods for estimating prediction error and providing statistical guarantees [15, 16, 17]. While this method has been utilized across disciplines for decades, its biggest limitation is the availability of data. The holdout method requires a validation set that is only used to assess the performance of the prediction model. If data is scarce, this can lead to test-set contamination or a small test-set size. However, in the context of data-driven reachability analysis, especially simulation-based approaches, generating new data sets avoids these issues and can be computationally efficient.
@@ II Problem Statement
To compute such an approximation, we first endow both $X_{0}$ and $D$ with probability distributions $\mu_{X_{0}}$ and $\mu_{D}$, respectively. From these distributions, we then draw samples $\delta^{(i)}=\Phi(t_{1};t_{0},x_{0i},d_{i}),i=1,\dotsc,N$ where $x_{01},\dotsc,x_{0N}\overset{i.i.d}{\sim}\mu_{X_{0}}$, $d_{1},\dotsc,d_{N}\overset{i.i.d}{\sim}\mu_{D}$. We compute a reachable set estimate in the form of a sublevel set of a parameterized function
@@
Given the samples $\delta^{(1)},\dots,\delta^{(N)}$ and a desired confidence parameter $\beta$, we wish to find the minimum-volume reachable set that contains the samples and satisfies the probabilistic guarantee $\mathbf{P}\{V(\hat{R}(\theta))>\epsilon\}\leq\beta$. The violation probability $V(\hat{R}(\theta))$, or true error $e$, of the reachable set estimate, $\hat{R}(\theta)$, is defined as the probability that an unseen scenario will violate the bounds of the reachable set:
@@
In other words, if $V(\hat{R}(\theta))\leq\epsilon$, then our reachable set estimate is robust against constraint violation at level $\epsilon$. Since the true error is not an observable quantity, we utilize the empirical error. Given a sample set $\Delta_{S}$, the empirical error $\hat{e}$, or test error, is the observed number of scenarios that violate the boundary of the reachable set estimate:
$\hat{e}\equiv\hat{V}(\hat{R}(\theta))$ $\equiv\mathbf{P}_{\delta_{s}\sim\Delta_{s}}\{\delta\not\in\hat{R}(\theta)\}$ (3)
$\equiv\frac{1}{M}\sum^{M}_{i=1}\mathds{1}_{\hat{R}(\theta)}(\delta_{i})$ (4)
@@
$\mathds{1}_{\hat{R}}(\theta)(\delta_{i})=\Bigg{\{}\begin{array}[]{ll}1&\text{if }\delta_{i}\notin\hat{R}(\theta)\\ 0&\text{else}.\\ \end{array}$ (5)
@@ II-C Nonconvex Scenario Reachability Analysis
To calculate reachable set estimates of the from (1), we utilize scenario optimization. Scenario optimization is an approach to solving chance-constrained optimization problems by solving a non-probabilistic relaxation of the original problem [25]. We fix a functional ${\rm Vol}:\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}_{\geq 0}$ that acts as a proxy for the volume of $R(\theta)$. This motivates the following scenario program:
[…]
@@ II-D Wait-and-Judge
Given a scenario program of the form (6), one approach for computing probabilistic bounds is wait-and-judge [18]. The wait-and-judge perspective, combined with nonconvex scenario optimization [26, 27], determines $\epsilon$ a-posteriori as a function of support scenarios. A support scenario is any scenario whose removal changes the solution to (6). This approach requires re-solving the scenario program upon removal of each individual scenario. Therefore, this approach can be computationally intensive for systems with high dimensionality or complexity. Since we calculate the probability measure of our reachable set after solving the optimization problem, the nonconvex scenario approach is amenable to alternative statistical analysis techniques, such as the holdout method, as discussed in Section III.
@@
To calculate probabilistic bounds for (6), we employ the holdout method given a sample set $\Delta_{S}$ of size $M$. We would like to calculate the probability of observing at most $k$ reachable boundary violations out of $M$ samples. From a statistical perspective, this can be viewed as computing a tail bound for a binomial distribution [28]. To obtain the largest true error such that the probability of $\frac{k}{M}$ or more samples violating the reachable set boundary is at least $\beta$, we use a binomial tail inversion. This calculates the bound on the true error given the empirical error $\hat{e}$, or the empirical count of boundary violations $\hat{k}$, and confidence $\beta$.
@@
(Binomial Tail Inversion) For $k$ violations out of $M$ newly sampled scenarios and all $\beta\in(0,1]$:
$\overline{\text{Bin}}(k,M,\beta)=\max_{e}\Bigl{\{}e:\text{Bin}\Bigl{(}k,M,e\Bigr{)}\geq\beta\Bigr{\}},\,\,\text{where}$ (7)
$\text{Bin}\Bigl{(}k,M,e\Bigr{)}=\sum_{j=0}^{k}\binom{M}{j}e^{j}(1-e)^{M-j}.$ (8)
@@ III The Holdout Method
The holdout method is a foundational paradigm of data-driven methodologies. It employs a test set of $M$ fresh scenarios to provide a good risk estimate for a given model [29]. Therefore, we draw a new set of samples $\delta^{(i)}_{s}=\Phi(t_{1};t_{0},x_{0i}^{s},d_{i}^{s}),i=1,\dotsc,M$ where $x_{01}^{s},\dotsc,x_{0M}^{s}\overset{i.i.d}{\sim}\mu_{X_{0}}$, $d_{1}^{s},\dotsc,d_{M}^{s}\overset{i.i.d}{\sim}\mu_{D}$, and test the accuracy of our estimate $\hat{R}(\theta)$ on $\delta_{s}^{(1)},\dotsc,\delta_{s}^{(M)}$. Given $k$ reachable set violations out of $M$ samples, or the empirical error (4), we are able to use a binomial tail bound, as introduced in Section II-E, to calculate a bound on the true error (2) of the reachable set. We formulate this in the following theorem:
@@
Computation. To calculate the binomial tail inversion (cf. Definition 1), we use the SciPy Python library to solve an optimization problem with the sequence’s cumulative distribution function (CDF) as a constraint, as seen in (7). SciPy offers both an optimization framework and a statistics library with built-in functions for computing binomial coefficients, such as the CDF. Since the binomial tail inversion reduces to inverting a one-dimensional function, it can be computed efficiently using standard optimization routines such as Quasi-Newton methods. Additionally, one can exploit the monotonically decreasing property of the function $e\mapsto\mathrm{Bin}(k,M,e)$ on the interval $[0,1]$ and utilize a simple bisection method.
Order-wise scaling of binomial tail inversion. In general, $\overline{\text{Bin}}(k,M,\beta)$ does not have a closed-form solution. However, we can readily obtain an order-wise scaling to illustrate the dependence on $k$, $M$, and $\beta$. First, suppose that no violations were observed on our holdout dataset ($k=0$). Then, we have the following bound $\overline{\text{Bin}}(0,M,\beta)\leq\frac{\log(1/\beta)}{M}$ [28, Corollary 3.4]. This illustrates a $1/M$ dependence (known as a “fast-rate” in the statistics literature) in the number of holdout examples $M$, in addition to a logarithmic dependence, $\log(\frac{1}{\beta})$, on the failure probability $\beta$. In the general case, when $k>0$, we have the looser scaling
$\overline{\text{Bin}}(k,M,\beta)\leq\frac{k}{M}+O\Bigl{(}\sqrt{\frac{\log(\frac{1}{\beta})}{M}}\Bigr{)},$ (10)
which is the typical scaling in $M$ predicted from the Central Limit Theorem.
Zero violation reachable sets. One drawback of the holdout method, compared to approaches such as wait-and-judge which do not use a separate holdout dataset, is that it is not possible to guarantee that the number of violations calculated from the holdout set is zero, even if a feasible solution to (6) is found on the training set. However, we show in Section IV, that while the wait-and-judge method computes a reachable set estimate with zero training violations, the final violation probability, $V(\hat{R}(\theta))$, is substantially more conservative than the holdout method.
Marginal probability over training and holdout data. Theorem 1 provides a bound over the probability of the $M$ holdout examples $\hat{\Delta}_{s}=\{\delta_{s}^{(i)}\}_{i=1}^{M}$. On the other hand, the reachable set $\hat{R}(\theta)$ is calculated as a function of the $N$ training examples $\hat{\Delta}=\{\delta^{(i)}\}_{i=1}^{N}$. It is also possible to obtain a bound which holds over the probability of all $N+M$ data points, which is more comparable to the guarantees provided by scenario optimization. In fact, the same bound from Theorem 1 holds by the tower property:
$\mathbf{P}_{(\hat{\Delta},\hat{\Delta}_{s})}\{V(\hat{R}(\theta))>\overline{\text{Bin}}(\hat{k},M,\beta)\}\leq\beta.$
@@ IV-A Reachable Sets
Observe that (11) corresponds to the program (6), where we take
@@
Further, we let Vol$(\hat{R}(\theta))=\sqrt{\sum_{i=1}^{m}\sigma_{i}^{2}}$ be a proxy for the volume of $\hat{R}(\theta)$. To solve (11), we first set the initial centers, $\mu_{i}$, of the RBFs to be close to optimal using a $k$-means clustering algorithm and choose the initial widths, $\sigma_{i}$, of our RBFs arbitrarily. We then use the SciPy Python library to solve this optimization problem using Sequential Least Squares Programming.
We take $\gamma=0.25$ to be the threshold of our RBF and consider 3000 samples. To apply the holdout method, we partition our data set into a training and test set, $N+M=3000$, and take $\beta=10^{-9}$. We compute reachable set estimates and $\epsilon$ a-posteriori for various combinations of $N$ and $M$, as seen in Tables I and II. Additionally, we investigate the difference in the volume proxy of our reachable set estimates, to gauge the effect of training set sample complexity. For a comparative baseline, we examine the wait-and-judge technique over $N=3000$ scenarios.
When we present the runtime of the holdout method, this measure includes the calculation of the reachable set over $N$ training samples, test-set generation of $M$ samples, and a-posteriori bound computation. In contrast, runtimes of the wait-and-judge method account for calculation of the reachable set over $N=3000$ samples and a-posteriori bound computation using support scenarios.
@@ IV-A1 Duffing Oscillator
The first example is a reachable set estimation problem for the nonlinear, time-varying system with dynamics: $\ddot{x}=-\alpha y+x-x^{3}+\gamma\cos(\omega t)$, with states $x,y\in\mathbb{R}$ and parameters $\alpha,\gamma,\omega\in\mathbb{R}$. This system is known as the Duffing oscillator, a nonlinear oscillator which exhibits chaotic behavior for certain values of $\alpha,\gamma$ and $\omega$, for instance $\alpha=0.05,\gamma=0.4,\omega=1.3$. The set of initial states is the interval such that $x(0)\in[0.95,1.05]$, $y(0)\in[-0.05,0.05]$, and we take $\mu_{X_{0}}$ to be the uniform random variable over this interval. The time range is $[t_{0},t_{1}]=[0,100]$.
@@
Holdout Method: Training ($N$) Testing ($M$) Vol$(\hat{R}(\theta))$ $\epsilon$
@@
We calculate a reachable set estimate, using two radial basis functions, such that $m=2$. We apply the holdout method, for various combinations of $N$ and $M$, all of which exhibited a runtime of approximately 10-15 sec. When varying the size of the training set, we encounter a significant decrease in volume of our reachable set estimates when the training set has 1000 or less samples. Further, extreme values of $N$, such as $N=10$ or $N=2990$, provide similarly poor probability measures. As seen in Fig. 2, it is best to balance the size of $N$ and $M$. We observe $N=1500$, $M=1500$ providing the smallest epsilon, $\epsilon=0.018$. In contrast, the wait-and-judge approach exhibited a runtime of approximately 22 min and resulted in $\epsilon=0.035$, with a volume proxy of Vol$(\hat{R}(\theta))=1.55$.
@@ IV-A2 Quadrotor
The next example is a reachable set estimation problem for a nonlinear model of a quadrotor used as an example in [30] and [31]. The dynamics for this system are
@@
$x(0)\in[-1.7,1.7],h(0)\in[0.3,2.0],\theta(0)\in[-\pi/12,\pi/12],$
$\dot{x}(0)\in[-0.8,0.8],\dot{h}(0)\in[-1.0,1.0],\dot{\theta}(0)\in[-\pi/2,\pi/2],$
the set of inputs is the set of constant functions $u_{1}(t)=u_{1}$, $u_{2}(t)=u_{2}$ $\forall t\in[t_{0},t_{1}]$, whose values lie in the interval $u_{1}\in[-1.5+g/K,1.5+g/K],u_{2}\in[-\pi/4,\pi/4]$, and we take $\mu_{X_{0}}$ and $\mu_{D}$ to be the uniform random variables defined over these intervals. The time range is $[t_{0},t_{1}]=[0,5]$.
We calculate a reachable set estimate, using three radial basis functions, such that $m=3$. We apply the holdout method, for various combinations of $N$ and $M$, all of which exhibited a runtime of approximately 35-50 sec. When varying the size of the training set, we observed very similar behavior as that depicted in Fig. 2 for the duffing oscillator. Extreme values of $N$ provide poor probability measures and less than 1000 training samples results in a decrease of the volume proxy of the reachable set estimate. In contrast, the wait-and-judge approach exhibited a runtime of approximately 5.5 hrs and resulted in $\epsilon=0.051$, with a volume proxy of Vol$(\hat{R}(\theta))=27.80$.
@@
Holdout Method: Training ($N$) Testing ($M$) Vol$(\hat{R}(\theta))$ $\epsilon$

### v2

@@ Abstract
©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Reachability analysis plays an integral role in analyzing system safety by determining the set of states a system can transition to in a finite time horizon [1]. Many set-based techniques [2, 3, 4, 5] have been developed to compute under- or over-approximated reachable sets to capture all possible trajectories and present guarantees on state-space safety. However, when there is limited information on a system’s dynamics, through only simulation or experimentation, we must estimate reachable sets and derive probabilistic guarantees of correctness directly from data. To provide probability measures for learned reachable sets, many approaches require a minimum number of samples [6, 7, 8, 9] or assumptions on the system dynamics [10, 11, 12], and are often computationally intensive [13, 14]. We advocate for the use of a general purpose method, the holdout method, that circumvents these limitations and improves the sample complexity and sharpness of existing probabilistic reachability bounds.
The holdout method, or cross-validation, is one of the simplest and most widely used methods for estimating prediction error and providing statistical guarantees [15, 16, 17, 18]. While this method has been utilized across disciplines for decades, its biggest limitation is the availability of data. The holdout method requires a validation set that is only used to assess the performance of the prediction model. If data is scarce, this can lead to test-set contamination or a small test-set size. However, in the context of data-driven reachability analysis, especially simulation-based approaches, generating new data sets avoids these issues and can be computationally efficient.
@@ II Problem Statement
To compute such an approximation, we first endow both $X_{0}$ and $D$ with probability distributions $\mu_{X_{0}}$ and $\mu_{D}$, respectively. Let $\Delta$ be the resulting probability space from which we draw samples $\delta^{(i)}=\Phi(t_{1};t_{0},x_{0i},d_{i}),i=1,\dotsc,N$ where $x_{01},\dotsc,x_{0N}\overset{i.i.d}{\sim}\mu_{X_{0}}$, $d_{1},\dotsc,d_{N}\overset{i.i.d}{\sim}\mu_{D}$. We compute a reachable set estimate in the form of a sublevel set of a parameterized function
@@
Given the samples $\delta^{(1)},\dots,\delta^{(N)}$ and a desired confidence parameter $\beta$, we wish to find the minimum-volume reachable set that contains the samples and satisfies the probabilistic guarantee $\mathbf{P}\{V(\hat{R}(\theta))>\epsilon\}\leq\beta$. The violation probability $V(\hat{R}(\theta))$, or true error $e$, of the reachable set estimate, $\hat{R}(\theta)$, is defined as the probability that an unseen scenario will violate the reachable set:
@@
In other words, if $V(\hat{R}(\theta))\leq\epsilon$, then our reachable set estimate is robust against constraint violation at level $\epsilon$. Since the true error is not an observable quantity, we utilize the empirical error. Given a new sample set of size $M$ drawn from $\Delta$, the empirical error $\hat{e}$, or test error, is the observed number of scenarios that violate the reachable set estimate:
$\hat{e}\equiv\hat{V}(\hat{R}(\theta))$ $\equiv\mathbf{P}_{\delta_{s}\sim\Delta}\{\delta_{s}\not\in\hat{R}(\theta)\}$ (3)
$\equiv\frac{1}{M}\sum^{M}_{i=1}\mathds{1}_{\hat{R}(\theta)}(\delta_{s}^{(i)})$ (4)
@@
[…]
@@ II-C Nonconvex Scenario Reachability Analysis
To calculate reachable set estimates of the form (1), we utilize scenario optimization. Scenario optimization is an approach to solving chance-constrained optimization problems by solving a non-probabilistic relaxation of the original problem [26]. We fix a functional ${\rm Vol}:\mathbb{R}^{n_{\theta}}\rightarrow\mathbb{R}_{\geq 0}$ that acts as a proxy for the volume of $R(\theta)$. This motivates the following scenario program:
$\underset{\theta}{\text{minimize}}$ $\rm{Vol}(\theta)$ (6)
@@ II-D Wait-and-Judge
Given a scenario program of the form (6), one approach for computing probabilistic bounds is wait-and-judge [19]. The wait-and-judge perspective, combined with nonconvex scenario optimization [27], determines $\epsilon$ a-posteriori as a function of support scenarios. A support scenario is any scenario whose removal changes the solution to (6). This approach requires re-solving the scenario program upon removal of each individual scenario. Therefore, this approach can be computationally intensive for systems with high dimensionality or complexity. We refer readers to [19] for more details on computing $\epsilon$. Since we calculate the probability measure of our reachable set after solving the optimization problem, the nonconvex scenario approach is amenable to alternative statistical analysis techniques, such as the holdout method, as discussed in Section III.
@@
To calculate probabilistic bounds for (6), we employ the holdout method given a sample set of size $M$. We would like to calculate the probability of observing at most $k$ reachable set violations out of $M$ samples. From a statistical perspective, this can be viewed as computing a tail bound for a binomial distribution. We use a binomial tail inversion to obtain the largest true error $e$ such that the probability of $k$ or more samples violating the reachable set is at least $\beta$ [28]. This calculates the bound on the true error given the empirical error $\hat{e}$, or the empirical count of boundary violations $\hat{k}$, and confidence $\beta$.
@@
(Binomial Tail Inversion) For $\hat{k}$ violations out of $M$ newly sampled scenarios and all $\beta\in(0,1]$:
$\overline{\text{Bin}}(\hat{k},M,\beta)=\max_{e}\Bigl{\{}e:\text{Bin}\Bigl{(}\hat{k},M,e\Bigr{)}\geq\beta\Bigr{\}},\,\,\text{where}$ (7)
$\text{Bin}\Bigl{(}\hat{k},M,e\Bigr{)}=\sum_{j=0}^{\hat{k}}\binom{M}{j}e^{j}(1-e)^{M-j}.$ (8)
@@ III The Holdout Method
The holdout method is a foundational paradigm of data-driven methodologies. It employs a test set of $M$ fresh scenarios to provide a good risk estimate for a given model [29]. Therefore, we draw a new set of samples $\delta^{(i)}_{s}=\Phi(t_{1};t_{0},x_{0i}^{s},d_{i}^{s}),i=1,\dotsc,M$ where $x_{01}^{s},\dotsc,x_{0M}^{s}\overset{i.i.d}{\sim}\mu_{X_{0}}$, $d_{1}^{s},\dotsc,d_{M}^{s}\overset{i.i.d}{\sim}\mu_{D}$, and test the accuracy of our estimate $\hat{R}(\theta)$ on $\delta_{s}^{(1)},\dotsc,\delta_{s}^{(M)}$. Given $\hat{k}$ reachable set violations out of $M$ samples, or the empirical error (4), we are able to use a binomial tail bound, as introduced in Section II-E, to calculate a bound on the true error (2) of the reachable set. We formulate this in the following theorem:
@@
Computation. To calculate the binomial tail inversion (cf. Definition 1), we use the SciPy Python library to solve an optimization problem with the sequence’s cumulative distribution function (CDF) as a constraint, as seen in (7). Since the binomial tail inversion reduces to inverting a one-dimensional function, it can be computed efficiently using standard optimization routines, such as Quasi-Newton methods. Additionally, one can exploit the monotonically decreasing property of the function $e\mapsto\mathrm{Bin}(k,M,e)$ on the interval $[0,1]$ and utilize a simple bisection method.
Order-wise scaling of binomial tail inversion. In general, $\overline{\text{Bin}}(\hat{k},M,\beta)$ does not have a closed-form solution. However, we can obtain an order-wise scaling to illustrate the dependence on $\hat{k}$, $M$, and $\beta$. First, suppose that no violations were observed on our holdout dataset ($\hat{k}=0$). Then, we have the following bound $\overline{\text{Bin}}(0,M,\beta)\leq\frac{\log(1/\beta)}{M}$ [28, Corollary 3.4]. This illustrates a $1/M$ dependence (known as a “fast-rate” in statistics) in the number of holdout samples $M$, in addition to a logarithmic dependence, $\log(\frac{1}{\beta})$, on the failure probability $\beta$. In the general case, when $\hat{k}>0$, we have the looser scaling
$\overline{\text{Bin}}(\hat{k},M,\beta)\leq\frac{\hat{k}}{M}+O\Bigl{(}\sqrt{\frac{\log(\frac{1}{\beta})}{M}}\Bigr{)},$ (10)
which is the typical scaling in $M$ predicted by the Central Limit Theorem.
Zero violation reachable sets. The holdout method, compared to approaches, such as wait-and-judge (which do not require a separate holdout dataset), is not able to guarantee zero violations on the complete dataset. However, we show in Section IV, that while the wait-and-judge method computes a reachable set estimate with zero dataset violations, the final violation probability, $V(\hat{R}(\theta))$, is substantially more conservative than the holdout method.
Marginal probability over training and holdout data. Theorem 1 provides a bound over the probability of the $M$ holdout samples $\{\delta_{s}^{(i)}\}_{i=1}^{M}$. On the other hand, the reachable set $\hat{R}(\theta)$ is calculated as a function of the $N$ training samples $\{\delta^{(i)}\}_{i=1}^{N}$. It is also possible to obtain a bound which holds over the probability of all $N+M$ data points, which is more comparable to the guarantees provided by scenario optimization. In fact, the same bound from Theorem 1 holds by the tower property:
$\mathbf{P}_{(\{\delta^{(i)}\}_{i=1}^{N},\{\delta_{s}^{(i)}\}_{i=1}^{M})}\{V(\hat{R}(\theta))>\overline{\text{Bin}}(\hat{k},M,\beta)\}\leq\beta.$
@@ IV-A Reachable Sets
Observe that (11) corresponds to the program (6), where
@@
[…]
We take $\gamma=0.25$ as the threshold of our RBF and consider 3000 samples. To apply the holdout method, we partition our data set into a training and test set, $N+M=3000$, and take $\beta=10^{-9}$. We compute reachable set estimates and $\epsilon$ a-posteriori for various combinations of $N$ and $M$, as seen in Tables I and II. Additionally, we investigate the difference in volume proxy of our reachable set estimates, to gauge the effect of training set sample complexity. For a comparative baseline, we examine the wait-and-judge technique over $N=3000$ scenarios. When presenting the runtime of the holdout method, this measure includes the calculation of the reachable set over $N$ training samples, test-set generation of $M$ samples, and a-posteriori bound computation. In contrast, runtimes of the wait-and-judge method account for calculation of the reachable set over $N=3000$ samples and a-posteriori bound computation using support scenarios.
@@ IV-A1 Duffing Oscillator
The first example is a reachable set estimation problem for the nonlinear, time-varying system with dynamics: $\ddot{x}=-\alpha y+x-x^{3}+\gamma\cos(\omega t)$, with states $x,y\in\mathbb{R}$ and parameters $\alpha,\gamma,\omega\in\mathbb{R}$. This system is known as the Duffing oscillator, a nonlinear oscillator which exhibits chaotic behavior for certain values of $\alpha,\gamma$ and $\omega$, for instance $\alpha=0.05,\gamma=0.4,\omega=1.3$. The set of initial states is the interval such that $x(0)\in[0.95,1.05]$, $y(0)\in[-0.05,0.05]$, and we take $\mu_{X_{0}}$ to be the uniform random variable over this interval. The time range is $t_{0},t_{1}]=[0,100]$.
@@
[…]
@@
[…]
@@ IV-A2 Quadrotor
The next example is for a nonlinear model of a quadrotor used as an example in [30, 14, 31]. The dynamics for this system are
@@
[…]
the set of inputs is the set of constant functions $u_{1}(t)=u_{1}$, $u_{2}(t)=u_{2}$ $\forall t\in[t_{0},t_{1}]$, whose values lie in the interval $u_{1}\in[-1.5+g/K,1.5+g/K],u_{2}\in[-\pi/4,\pi/4]$, and we take $\mu_{X_{0}}$ and $\mu_{D}$ to be the uniform random variables defined over these intervals. The time range is $t_{0},t_{1}]=[0,5]$.
[…]
Sample Complexity. In the section above, we utilize equal size datasets to compare the wait-and-judge and holdout method, highlighting the improved accuracy and computational cost of the holdout method. To demonstrate futher improvements in sample complexity, let us examine the duffing oscillator example with sample size $N+M=2000$ where $N=1000$ and $M=1000$. We apply the holdout method and calculate $\epsilon=0.0263$, achieving a $1\%$ increase in accuracy with $1000$ fewer samples when compared to the wait-and-judge approach on $3000$ samples.
@@
[…]